A 249−mL benzene solution containing 2.35 g of an organic polymer has an osmotic pressure of 7.91 mm Hg at 19°C. Calculate the molar mass of the polymer.
Answer #1

We Have

Osmotic Pressure P = 7.91 mm Hg = 7.91/760 = 0.0104 atm

Temperature T = 190 C = 19 + 273 = 292 K

Volume = 249 mL = 0.249 L

mass of polymer = 2.35 g

R = Ideal Gas Constant = 0.082057 L atm mol-1 K-1

C = Molar Concentration = ?

So We know the formula

Osmotic Pressure = CRT

So

0.0104 atm = C 0.082057 L atm mol-1 K-1  292 K

C = 0.0104 atm/ 23.96 atm mol-1 = 4.34 10-4 M

Now Lets Calculate the number of moles by given volume

Moles of Polymer = 4.34 10-4 M   0.249 L = 1.08   10-4 mol

From this we can get Molar mass now

Molar mass of Polymer = mass of polymer / moles of Polymer = 2.35 g / 1.08   10-4 mol

= 2.17   104 g/mol

So Our Answer is = 2.17 104 g/mol